OpenStudy (anonymous):

Evaluate the function T(x) = 19 − x, when x = −3 (Points : 1) 16 57 22 −16

OpenStudy (anonymous):

T(-3)=19-(-3)=19+3=22

OpenStudy (anonymous):

r u sure

OpenStudy (anonymous):

yup :)